# Risk premium and rough volatility

### v1

@@ 1. Rough volatility models and change of measure
(1.12) $\widetilde{S}_{0}-\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}]=\lim_{n\uparrow\infty}\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{\iota_{n}}1{1}_{\{T>\iota_{n}\}}].$
Girsanov’s Theorem further gives $\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}1{1}_{\{T>\iota_{n}\}}]=\widetilde{S}_{0}\widetilde{\mathbb{P}}_{n}(T>\iota_{n})$, where ~$\widetilde{\mathbb{P}}_{n}$ is such that
(1.13) $\widetilde{W}^{n}_{t}=W^{\mathbb{Q}}_{t}-\int_{0}^{t\wedge\iota_{n}}v_{s}\mathrm{d}s$
is a ~$\widetilde{\mathbb{P}}_{n}$-Brownian motion. Note that, for $t<\iota_{n}$, $\widehat{Y}_{t}=\widetilde{Y}_{t}+\rho\int_{0}^{t}k(t,s)v_{s}\mathrm{d}s$, where $\widetilde{Y}_{t}=\int_{0}^{t}k(t,s)\mathrm{d}\widetilde{Z}^{n}_{s}$ and $\widetilde{Z}^{n}_{t}:=Z^{\mathbb{Q}}_{t}-\rho\int_{0}^{t}k(t,s)v_{s}\mathrm{d}s$ is also a ~$\widetilde{\mathbb{P}}_{n}$-Brownian motion. We conclude that, if $\rho\leq 0$, then $\widehat{Y}_{t}\geq\widetilde{Y}_{t}$ and
@@
Discussion
Under Assumption 1.1, consider $\rho\leq 0$, some valid function $\psi$ and kernel $k$, and the constant values $\gamma_{s}=\overline{\gamma}$ and $\overline{\mu}=\mu_{s}\leq r_{s}=\overline{r}$ for $\overline{\gamma},\overline{\mu},\overline{r}\in\mathbb{R}$ ensuring Assumption 1.1(iv)., so that Theorem 1.5 applies. In this scenario, a sufficient condition for the change of measure to be well defined is that the physical drift must be smaller than the risk-free rate.
1.1. Characterisation of rough volatility models via Generalised Fractional Operators
It is natural to represent rough volatility models in terms of fractional operators. In this section we present the Generalised Fractional Operators (GFO) and a representation result for rough volatility models in terms of GFO. To do so, let us first present the GFO introduced in [15, Definition 1.1] defined as follows.
@@ 1.1. Characterisation of rough volatility models via Generalised Fractional Operators
(1.14) $(\mathcal{G}^{\alpha}f)(t):=\left\{\begin{array}[]{ll}\displaystyle\int_{0}^{t}(f(s)-f(0))\frac{\mathrm{d}}{\mathrm{d}t}k(t-s)\mathrm{d}s,&\text{if }\alpha\in[0,1-\beta),\\ \displaystyle\frac{\mathrm{d}}{\mathrm{d}t}\int_{0}^{t}(f(s)-f(0))k(t-s)\mathrm{d}s,&\text{if }\alpha\in(-\beta,0).\end{array}\right.$
@@
(1.15) $\mathtt{k}_{\alpha}(u):=u^{\alpha}1{1}_{\{u\geq 0\}},$
as well as the set
(1.16) $\Lambda_{\beta,H}:=\Big{\{}\lambda\in\mathcal{C}^{\beta}\text{ for some $\beta\in(0,1]$ such that $H_{-}\in(-\beta,0)$ and $\lambda_{0}=0$}\Big{\}}.$
To this particular power-law kernel, the GFO (from Definition 1.6, since $H_{-}\in(-\frac{1}{2},0)$) reads
@@
From the properties of GFO [15, Proposition 1.2], then $\mathfrak{G}^{+}\lambda\in\mathcal{C}^{\beta+H_{+}}$ as soon as $\lambda\in\Lambda_{\beta,H}$.
Corollary 1.7 (GFO representation of rough volatility).
@@
The fact that $\int_{0}^{\cdot}\mathtt{k}_{H_{-}}(\cdot-s)\mathrm{d}Z^{\mathbb{Q}}_{s}=\mathfrak{G}^{-}Z^{\mathbb{Q}}$ is straightforward by the properties of GFO in [15, Proposition 1.4]. Furthermore, for any $\lambda\in\Lambda_{\beta,H}$ and any $t\in\mathbb{T}$,
@@ 2. Modelling the risk premium process: A practical approach
In practice, the process $\lambda$ is directly modelled without resorting to a change of measure starting from $\gamma$. We now consider different modelling choices for the risk premium $\lambda$ and analyse some of its properties. In spite of the formal derivation of Theorem 1.5, a numerical treatment of the integral $\int_{0}^{t}\bullet\;\mathrm{d}s$ is rather intricate. To overcome this issue, Bayer, Friz and Gatheral [3] elegantly came up with the forward variance form of rough volatility in the spirit of Bergomi [5]. We shall restrict ourselves to this functional form (defined in the following proposition) for the reminder of the section. Consider (1.5) with $\psi(t,x)=\xi_{0}(t)\mathrm{e}^{\nu x}$, with $\xi_{0}(t):=\mathbb{E}[v_{t}|\mathcal{F}_{0}]$ and $\nu>0$. Then the risk-neutral dynamics in forward variance form read
@@ 2.1. Risk premium driven by Itô diffusion
For $H\in(0,\frac{1}{2})$ and $\alpha\in\left(-\frac{1}{2},0\right)$, consider $\lambda=\mathfrak{b}\mathcal{G}^{\alpha}Y^{\mathbb{Q}}\in\mathcal{C}^{\alpha+\frac{1}{2}}$, where $\mathfrak{b}:=\mathfrak{B}(H_{+},\alpha+1)^{-1}$ and
$Y^{\mathbb{Q}}_{t}=\int_{0}^{t}b(s,Y^{\mathbb{Q}}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,Y^{\mathbb{Q}}_{s})\mathrm{d}X^{\mathbb{Q}}_{s},$
where $b(\cdot)$ and $\sigma(\cdot,\cdot)$ satisfy Yamada-Watanabe conditions [20, Section 5.2, Proposition 2.13] for pathwise uniqueness ensuring that a weak solution exists. Then $\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\in\mathcal{C}^{H+\alpha+1}$ and
@@
(2.4) ×$\times\exp\left\{\frac{\nu^{2}}{2}\left(\frac{\mathtt{k}_{2H}(t-s)}{2H}+\frac{\mathtt{k}_{2(H+1)}(t-s)}{2H_{+}^{2}(H+1)}+\rho\frac{\mathtt{k}_{2H_{+}}(t-s)}{H_{+}^{2}}\right)\right\},$
@@
We first prove (2.2), which follows from [15, Proposition 1.2] and the identities highlighted above in Corollary 1.7. Indeed, in view of Corollary 1.7 we only need to show that
$\int_{0}^{t}\mathtt{k}_{H_{-}}(t-s)\lambda_{s}\mathrm{d}s=(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}})(t).$
Replacing the expression for $\lambda$ in the integral and using stochastic Fubini theorem, we obtain
@@
Now, direct computations for the inner integral yields
@@
×$\qquad\times\mathbb{E}^{\mathbb{Q}}_{s}\left[\exp\left\{\nu\left[\int_{s}^{t}\mathtt{k}_{H_{-}}(t-u)\mathrm{d}Z^{\mathbb{Q}}_{u}+\int_{s}^{t}\mathtt{k}_{\alpha+H_{+}}(t-u)\mathrm{d}X^{\mathbb{Q}}_{u}\right]\right\}\right]$
@@
×$\qquad\times\exp\left\{\frac{\nu^{2}}{2}\left(\int_{s}^{t}\mathtt{k}_{2H_{-}}(t-u)\mathrm{d}u+\int_{s}^{t}\frac{\mathtt{k}_{2H+1}(t-u)}{H_{+}^{2}}\mathrm{d}u+\rho\int_{s}^{t}\frac{\mathtt{k}_{2H}(t-u)}{H_{+}}\mathrm{d}u\right)\right\}$
@@
×$\qquad\times\exp\left\{\frac{\nu^{2}}{2}\left(\frac{\mathtt{k}_{2H}(t-s)}{2H}+\frac{\mathtt{k}_{2(H+1)}(t-s)}{2H_{+}^{2}(H+1)}+\rho\frac{\mathtt{k}_{2H_{+}}(t-s)}{H_{+}^{2}}\right)\right\}.$
@@
Since the instantaneous variance in this model is log-Normal, the results in [18, Proposition 3.1] and numerical methods therein still apply for the VIX with minimal changes.
@@ 2.2. A risk premium driven by a CIR process
×$\qquad\times\mathbb{E}\left[\exp\Big{\{}\nu\left(\left(\mathfrak{G}^{-}Z^{\mathbb{Q}}\right)(t)-\left(\mathfrak{G}^{-}Z^{\mathbb{Q}}\right)(u,t)\right)\Big{\}}\right]$
×$\qquad\times\mathbb{E}\left[\exp\Big{\{}\nu\left(\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(t)-\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(u,t)\right)\Big{\}}\right],$
@@
where
$\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(t)-\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(u,t)=\int_{u}^{t}\mathtt{k}_{H_{-}}(t,s)Y^{\mathbb{Q}}_{s}\mathrm{d}s.$
This is, in spirit, similar to computing a bond price in the CIR model. To do so, define
$B(t,T):=\mathbb{E}\left[\exp\left(\nu\int_{t}^{T}\mathtt{k}_{H_{-}}(T,s)Y^{\mathbb{Q}}_{s}\mathrm{d}s\right)\bigg{|}\mathcal{F}_{t}\right],$
@@
where $\mathcal{B}(y,t,T)$ is such that $\mathcal{B}(Y^{\mathbb{Q}}_{t},t,T)$ is a solution to the moment generating function. With an ansatz of the type $\mathcal{B}(y,t,T)=\exp\{-yC(t,T)-A(t,T)\}$, we have at $(y,t,T)$,
@@
The second term cancels when $A(t,T)=-\kappa\theta\int_{t}^{T}C(u,T)\mathrm{d}t$, and a Riccati equation remains:
$\nu\mathtt{k}_{H_{-}}(T,t)-\partial_{t}C(t,T)-\kappa C(t,T)+\frac{\sigma^{2}}{2}C^{2}(t,T)=0,$
with boundary conditions $A(T,T)=C(T,T)=0$. ∎
Already in the uncorrelated case the computation of $\mathbb{E}^{\mathbb{Q}}[v_{t}]$ becomes very costly, having to solve a PDE for each time $t$. In the correlated case there is no hope to obtain any semi-analytic result since one would need to compute cross terms and there is no tool coming from Itô’s calculus available in that case.
@@ 3. Roughly extracting the Risk Premium from the Market
In this section we consider the risk premium process $\lambda$ to be deterministic, with the aim of obtaining a formula that links $\mathbb{P}$ and $\mathbb{Q}$ market observable quantities. The following theorem shows how to infer the risk premium from the market using forecasts under the physical measure and Variance Swap prices in the pricing measure.
@@
If $\mu=r$ almost surely, the Radon-Nikodym derivative (1.4) in Theorem 1.5 reads
(3.2) $\mathcal{D}^{\gamma}=\mathcal{E}\left(\int_{0}^{\cdot}\gamma_{s}\mathrm{d}W^{\mathbb{P}\perp}_{s}\right),$
where we recall $\lambda_{s}=\overline{\rho}\gamma_{s}$, and the inverse Radon-Nikodym derivative is given by
[…]
Then, the conditional change of measure formula yields
@@
(3.5) $\mathbb{E}^{\mathbb{Q}}\left[v_{t}\reflectbox{$\mathcal{D}$}^{\gamma}_{t}|\mathcal{F}_{s}\right]$ $=\mathbb{E}^{\mathbb{Q}}\left[\left.\exp\left\{\nu\left(\int_{0}^{t}\mathtt{k}_{H_{-}}(t,s)\mathrm{d}Z^{\mathbb{Q}}_{s}+\int_{0}^{t}\lambda_{s}\mathtt{k}_{H_{-}}(t,s)\mathrm{d}s\right)\right\}\mathrm{e}^{-\int_{0}^{t}\gamma_{s}\mathrm{d}W^{\mathbb{Q}\perp}_{s}-\frac{1}{2}\int_{0}^{t}\gamma_{s}^{2}\mathrm{d}s}\right|\mathcal{F}_{s}\right]$
(3.6) $=\mathrm{e}^{\nu\int_{0}^{t}\lambda_{s}\mathtt{k}_{H_{-}}(t,s)\mathrm{d}s-\frac{1}{2}\int_{0}^{t}\gamma_{s}^{2}\mathrm{d}s}\mathbb{E}^{\mathbb{Q}}\left[\left.\exp\left\{\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,s)\mathrm{d}Z^{\mathbb{Q}}_{s}-\int_{0}^{t}\gamma_{s}\mathrm{d}W^{\mathbb{Q}\perp}_{s}\right\}\right|\mathcal{F}_{s}\right],$
@@
(3.7) $\mathbb{E}^{\mathbb{Q}}\left[v_{t}\reflectbox{$\mathcal{D}$}^{\gamma}_{t}|\mathcal{F}_{s}\right]$ $=\exp\Bigg{\{}\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,s)\lambda_{s}\mathrm{d}s+\overbrace{\nu\int_{0}^{s}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}Z^{\mathbb{Q}}_{u}-\int_{0}^{s}\gamma_{u}\mathrm{d}W^{\mathbb{Q}\perp}_{u}}^{\mu}$
$\qquad+\frac{1}{2}\Bigg{(}\underbrace{\nu^{2}\int_{s}^{t}k^{2}(t,u)\mathrm{d}u+\int_{s}^{t}\gamma_{u}^{2}\mathrm{d}u-2\nu\overline{\rho}\int_{s}^{t}\mathtt{k}_{H_{-}}(t,u)\gamma_{u}\mathrm{d}u}_{\sigma^{2}}-\int_{0}^{t}\gamma_{s}^{2}\mathrm{d}s\Bigg{)}\Bigg{\}}$
@@
$\qquad\underbrace{+\exp\Bigg{\{}\nu\int_{0}^{s}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}Z^{\mathbb{Q}}_{u}+\frac{\nu^{2}}{2}\int_{s}^{t}\mathtt{k}_{H_{-}}^{2}(t,u)\mathrm{d}u+\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,s)\lambda_{s}\mathrm{d}s\Bigg{\}}}_{\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{s}]},$

### v2

@@ 1. Rough volatility models and change of measure
(1.11) $\widetilde{S}_{0}-\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}]=\lim_{n\uparrow\infty}\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{\iota_{n}}1\hskip-5.97508pt{1}_{\{T>\iota_{n}\}}].$
Girsanov’s Theorem further gives $\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}1\hskip-5.97508pt{1}_{\{T>\iota_{n}\}}]=\widetilde{S}_{0}\widetilde{\mathbb{P}}_{n}(T>\iota_{n})$, where $\widetilde{\mathbb{P}}_{n}$ is such that $\widetilde{W}^{n}_{t}=W^{\mathbb{Q}}_{t}-\int_{0}^{t\wedge\iota_{n}}v_{s}\mathrm{d}s$ is a $\widetilde{\mathbb{P}}_{n}$-Brownian motion. Note that, for $t<\iota_{n}$, $\widehat{Y}_{t}=\widetilde{Y}_{t}+\rho\int_{0}^{t}k(t,s)v_{s}\mathrm{d}s$, where $\widetilde{Y}_{t}=\int_{0}^{t}k(t,s)\mathrm{d}\widetilde{Z}^{n}_{s}$ and $\widetilde{Z}^{n}_{t}:=Z^{\mathbb{Q}}_{t}-\rho\int_{0}^{t}k(t,s)v_{s}\mathrm{d}s$ is also a $\widetilde{\mathbb{P}}_{n}$-Brownian motion. We conclude that, if $\rho\leq 0$, then $\widehat{Y}_{t}\geq\widetilde{Y}_{t}$ and
@@
Remark 1.7.
Under Assumption 1.3, consider $\rho\leq 0$, some valid function $\psi$ and kernel $k$, and the constant values $\gamma_{s}=\overline{\gamma}$ and $\overline{\mu}=\mu_{s}\leq r_{s}=\overline{r}$ for $\overline{\gamma},\overline{\mu},\overline{r}\in\mathbb{R}$ ensuring Assumption 1.3(iv), so that Theorem 1.6 applies. In this scenario, a sufficient condition for the change of measure to be well defined is that the physical drift must be smaller than the risk-free rate.
1.1. Rough volatility models via Generalised Fractional Operators
Many rough volatility models can be represented [horvath2024functional] in terms of Generalised Fractional Operators (GFOs), which are defined as follows [horvath2024functional, Definition 1.1]:
@@ 1.1. Rough volatility models via Generalised Fractional Operators
[…]
@@
(1.13) $\mathtt{k}_{\alpha}(u):=u^{\alpha}1\hskip-5.97508pt{1}_{\{u\geq 0\}},$
[…]
@@
From the properties of GFO [horvath2024functional, Proposition 1.2], then $\mathfrak{G}^{+}\lambda\in\mathcal{C}^{\beta+H_{+}}$ as soon as $\lambda\in\Lambda_{\beta,H}$.
Corollary 1.9 (GFO representation).
@@
The fact that $\int_{0}^{\cdot}\mathtt{k}_{H_{-}}(\cdot-s)\mathrm{d}Z^{\mathbb{Q}}_{s}=\mathfrak{G}^{-}Z^{\mathbb{Q}}$ is straightforward by the properties of GFO in [horvath2024functional, Proposition 1.4]. Furthermore, for any $\lambda\in\Lambda_{\beta,H}$ and any $t\in\mathbb{T}$,
@@ 2. Modelling the risk premium process: A practical approach
In practice, the process $\lambda$ is directly modelled without resorting to a change of measure starting from $\gamma$. We now consider different modelling choices for the risk premium $\lambda$ and analyse some of its properties. In spite of the formal derivation of Theorem 1.6, a numerical treatment of the integral $\int_{0}^{t}\bullet\;\mathrm{d}s$ is rather intricate. To overcome this issue, Bayer, Friz and Gatheral [BFG15] elegantly came up with the forward variance form of rough volatility in the spirit of Bergomi [Bergomi]. We shall restrict ourselves to this functional form (defined below in (2.1)) for the remainder of the section. Consider (1.7) with $\psi(t,x)=\xi_{0}(t)\mathrm{e}^{\nu x}$, with $\xi_{0}(t):=\mathbb{E}[v_{t}|\mathcal{F}_{0}]$ and $\nu>0$. Then the risk-neutral dynamics in forward variance form read
@@ 2.1. Risk premium driven by Itô diffusion
For $H\in(0,\frac{1}{2})$, $\alpha\in\left(-\frac{1}{2},0\right)$, let $\lambda:=\mathfrak{b}\mathcal{G}^{\alpha}Y^{\mathbb{Q}}\in\mathcal{C}^{\alpha+\frac{1}{2}}$ with $\mathfrak{b}:=\mathfrak{B}(H_{+},\alpha+1)^{-1}$ and $Y^{\mathbb{Q}}_{t}=\int_{0}^{t}b(s,Y^{\mathbb{Q}}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,Y^{\mathbb{Q}}_{s})\mathrm{d}X^{\mathbb{Q}}_{s}$, where $b(\cdot)$ and $\sigma(\cdot,\cdot)$ satisfy the Yamada-Watanabe conditions [karatzas2012brownian, Section 5.2, Proposition 2.13] for pathwise uniqueness ensuring a weak solution. Then $\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\in\mathcal{C}^{H+\alpha+1}$ and
@@
(2.4) $\times\exp\left\{\frac{\nu^{2}}{2}\left(\frac{\mathtt{k}_{2H}(t-s)}{2H}+\frac{\mathtt{k}_{2(H+1)}(t-s)}{2H_{+}^{2}(H+1)}+\rho\frac{\mathtt{k}_{2H_{+}}(t-s)}{H_{+}^{2}}\right)\right\},$
@@
We first prove (2.2), which follows from [horvath2024functional, Proposition 1.2] and the identities highlighted above in Corollary 1.9. Indeed, in view of Corollary 1.9 we only need to show that $\int_{0}^{t}\mathtt{k}_{H_{-}}(t-s)\lambda_{s}\mathrm{d}s=(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}})(t)$. Replacing the expression for $\lambda$ in the integral and using stochastic Fubini theorem, we obtain
@@
Now, direct computations for the inner integral yield
@@
$\qquad\times\mathbb{E}^{\mathbb{Q}}_{s}\left[\exp\left\{\nu\left[\int_{s}^{t}\mathtt{k}_{H_{-}}(t-u)\mathrm{d}Z^{\mathbb{Q}}_{u}+\int_{s}^{t}\mathtt{k}_{\alpha+H_{+}}(t-u)\mathrm{d}X^{\mathbb{Q}}_{u}\right]\right\}\right]$
@@
$\qquad\times\exp\left\{\frac{\nu^{2}}{2}\left(\int_{s}^{t}\mathtt{k}_{2H_{-}}(t-u)\mathrm{d}u+\int_{s}^{t}\frac{\mathtt{k}_{2H+1}(t-u)}{H_{+}^{2}}\mathrm{d}u+\rho\int_{s}^{t}\frac{\mathtt{k}_{2H}(t-u)}{H_{+}}\mathrm{d}u\right)\right\}$
@@
$\qquad\times\exp\left\{\frac{\nu^{2}}{2}\left(\frac{\mathtt{k}_{2H}(t-s)}{2H}+\frac{\mathtt{k}_{2(H+1)}(t-s)}{2H_{+}^{2}(H+1)}+\rho\frac{\mathtt{k}_{2H_{+}}(t-s)}{H_{+}^{2}}\right)\right\}.$
@@
Since the instantaneous variance in this model is log-Normal, the results in [jacquier2018vix, Proposition 3.1] and numerical methods therein still apply for the VIX with minimal changes.
@@ 2.2. A risk premium driven by a CIR process
$\qquad\times\mathbb{E}\left[\exp\Big\{\nu\left(\left(\mathfrak{G}^{-}Z^{\mathbb{Q}}\right)(t)-\left(\mathfrak{G}^{-}Z^{\mathbb{Q}}\right)(u,t)\right)\Big\}\right]$
$\qquad\times\mathbb{E}\left[\exp\Big\{\nu\left(\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(t)-\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(u,t)\right)\Big\}\right],$
@@
where $\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(t)-\left(\mathcal{G}^{\alpha+H_{+}}Y^{\mathbb{Q}}\right)(u,t)=\int_{u}^{t}\mathtt{k}_{H_{-}}(t,s)Y^{\mathbb{Q}}_{s}\mathrm{d}s$. This is, in spirit, similar to computing a bond price in the CIR model. To do so, define
(2.6) $B(t,T):=\mathbb{E}\left[\exp\left(\nu\int_{t}^{T}\mathtt{k}_{H_{-}}(T,s)Y^{\mathbb{Q}}_{s}\mathrm{d}s\right)\bigg|\mathcal{F}_{t}\right],$
@@
where $\mathcal{B}(y,t,T)$ is such that $\mathcal{B}(Y^{\mathbb{Q}}_{t},t,T)=B(t,T)$ given in (2.6). With an ansatz of the type $\mathcal{B}(y,t,T)=\exp\{-yC(t,T)-A(t,T)\}$, we have at $(y,t,T)$,
@@
The last term cancels for $A(t,T)=-\kappa\theta\int_{t}^{T}C(u,T)\mathrm{d}t$, and the Riccati equation $\nu\mathtt{k}_{H_{-}}(T,t)-\partial_{t}C(t,T)-\kappa C(t,T)+\frac{\sigma^{2}}{2}C^{2}(t,T)=0$ remains, with $A(T,T)=C(T,T)=0$. ∎
Already in the uncorrelated case the computation of $\mathbb{E}^{\mathbb{Q}}[v_{t}]$ becomes very costly, having to solve a quadratic ODE (with time-dependent coefficients) for each time $t$. In the correlated case there is no hope to obtain any semi-analytic result since one would need to compute cross terms and there is no tool coming from Itô’s calculus available in that case. The approach in this section was essentially top-down, meaning that we specified a form for the market price of volatility risk $\lambda$ and deduced the shape of the model with this specification. Unfortunately, our analysis showed that this may ultimately not be so successful as the final form of the model is rather intricate, probably too much so for practical purposes. Alternatively, one may first infer some shape of $\lambda$ from market data (short rate of interest, expected returns and instantaneous volatility) and then use it to price options under $\mathbb{Q}$.
@@ 3. Roughly extracting the Risk Premium from the Market
We now consider the risk premium process $\lambda$ deterministic, and obtain a formula linking $\mathbb{P}$ and $\mathbb{Q}$ market observable quantities. The following theorem shows how to infer the risk premium from the market using forecasts under $\mathbb{P}$ and Variance Swap prices under $\mathbb{Q}$.
@@
If $\mu=r$ almost surely, the Radon-Nikodym derivative (1.3) in Theorem 1.6 reads $\mathcal{D}^{\gamma}=\mathcal{E}\left(\int_{0}^{\cdot}\gamma_{s}\mathrm{d}W^{\mathbb{P}\perp}_{s}\right)$, with $\lambda_{s}=\overline{\rho}\gamma_{s}$, and the inverse Radon-Nikodym derivative is given by $\reflectbox{$\mathcal{D}$}^{\gamma}:=\frac{1}{\mathcal{D}^{\gamma}}=\mathcal{E}\left(-\int_{0}^{\cdot}\gamma_{s}\mathrm{d}W^{\mathbb{Q}\perp}_{s}\right)$. Then, the conditional change of measure formula yields
@@
(3.3) $\mathbb{E}^{\mathbb{Q}}\left[v_{t}\reflectbox{$\mathcal{D}$}^{\gamma}_{t}|\mathcal{F}_{s}\right]$ $=\mathbb{E}^{\mathbb{Q}}\left[\left.\exp\left\{\nu\left(\int_{0}^{t}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}Z^{\mathbb{Q}}_{u}+\int_{0}^{t}\lambda_{u}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}u\right)\right\}\mathrm{e}^{-\int_{0}^{t}\gamma_{u}\mathrm{d}W^{\mathbb{Q}\perp}_{u}-\frac{1}{2}\int_{0}^{t}\gamma_{u}^{2}\mathrm{d}u}\right|\mathcal{F}_{s}\right]$
(3.4) $=\mathrm{e}^{\nu\int_{0}^{t}\lambda_{u}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}u-\frac{1}{2}\int_{0}^{t}\gamma_{u}^{2}\mathrm{d}u}\mathbb{E}^{\mathbb{Q}}\left[\left.\exp\left\{\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,s)\mathrm{d}Z^{\mathbb{Q}}_{s}-\int_{0}^{t}\gamma_{s}\mathrm{d}W^{\mathbb{Q}\perp}_{s}\right\}\right|\mathcal{F}_{s}\right],$
@@
(3.5) $\mathbb{E}^{\mathbb{Q}}\left[v_{t}\reflectbox{$\mathcal{D}$}^{\gamma}_{t}|\mathcal{F}_{s}\right]\$$=\exp\Bigg\{\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,u)\lambda_{u}\mathrm{d}u+\overbrace{\nu\int_{0}^{s}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}Z^{\mathbb{Q}}_{u}-\int_{0}^{s}\gamma_{u}\mathrm{d}W^{\mathbb{Q}\perp}_{u}}^{\mu}$
$\qquad+\frac{1}{2}\Bigg(\underbrace{\nu^{2}\int_{s}^{t}k^{2}(t,u)\mathrm{d}u+\int_{s}^{t}\gamma_{u}^{2}\mathrm{d}u-2\nu\overline{\rho}\int_{s}^{t}\mathtt{k}_{H_{-}}(t,u)\gamma_{u}\mathrm{d}u}_{\sigma^{2}}-\int_{0}^{t}\gamma_{u}^{2}\mathrm{d}u\Bigg)\Bigg\}$
@@
$\qquad\underbrace{+\exp\Bigg\{\nu\int_{0}^{s}\mathtt{k}_{H_{-}}(t,u)\mathrm{d}Z^{\mathbb{Q}}_{u}+\frac{\nu^{2}}{2}\int_{s}^{t}\mathtt{k}_{H_{-}}^{2}(t,u)\mathrm{d}u+\nu\int_{0}^{t}\mathtt{k}_{H_{-}}(t,u)\lambda_{u}\mathrm{d}u\Bigg\}}_{\mathbb{E}^{\mathbb{Q}}[v_{t}|\mathcal{F}_{s}]},$